OpenStudy (anonymous):

Write the equation of the line that passes through (-3, 4) with a slope of 2 in point-slope form.

OpenStudy (anonymous):

A. y - 4 = 2(x + 3) B. y + 4 = 2(x - 3) C. y + 3 = 2(x - 4) D. y - 3 = 2(x + 4)

OpenStudy (firekat97):

do you know this equation? (y - y1) = m(x - x1)

OpenStudy (anonymous):

Ive seen it before yes. m=2

OpenStudy (anonymous):

need some help?

OpenStudy (anonymous):

y+4=2(x-3) I think

OpenStudy (anonymous):

So B?

OpenStudy (firekat97):

well for the right hand side you have to do m(x - x1) so 2(x - -3) so 2(x +3) but the right hand side is correct and you get the basic idea :)

OpenStudy (firekat97):

Just be weary of signs though

OpenStudy (anonymous):

Ohhh okay so A

OpenStudy (anonymous):

Thank you! Could you help me on another with 2 pairs of points

OpenStudy (anonymous):

@FireKat97

OpenStudy (firekat97):

sureee

OpenStudy (anonymous):

Thank you!

OpenStudy (firekat97):

do you have a specific question? or just generally?

OpenStudy (firekat97):

no problem :)